# Mind Your Format: Towards Consistent Evaluation of In-Context Learning Improvements

### v1

@@ Abstract
Large language models demonstrate a remarkable capability for learning to solve new tasks from a few examples. The prompt template, or the way the input examples are formatted to obtain the prompt, is an important yet often overlooked aspect of in-context learning. In this work, we conduct a comprehensive study of the template format’s influence on the in-context learning performance. We evaluate the impact of the prompt template across 19 models (from 770M to 70B parameters) and 4 standard classification datasets. We show that a poor choice of the template can reduce the performance of the strongest models and inference methods to a random guess level. More importantly, the best templates do not transfer between different setups and even between models of the same family. Our findings show that the currently prevalent approach to evaluation, which ignores template selection, may give misleading results due to different templates in different works. As a first step towards mitigating this issue, we propose Template Ensembles that aggregate model predictions across several templates. This simple test-time augmentation boosts average performance while being robust to the choice of random set of templates.
@@ 1 Introduction
These few-shot or in-context Brown et al. (2020) learning (ICL) abilities of large models are a subject of frequent study, as the primary factors behind them are not yet fully understood. For example, one line of work investigates in-context learning within different theoretical frameworks Xie et al. (2022); Garg et al. (2022); Akyürek et al. (2023). In addition, multiple publications study the importance of different prompt attributes, such as the order of input demonstrations (Lu et al., 2022a) and their labels (Min et al., 2022d).
@@
As shown in Zhao et al. (2021); Min et al. (2022a), the prompt format (i.e., a transformation from a set of examples to a natural language input) is also highly important. However, this aspect is often overlooked in most existing studies. Namely, works proposing modifications of ICL frequently present their results for a specific template without specifying the criteria guiding its selection. Furthermore, even when the results are averaged over a set of templates, they are compared to methods that were evaluated on a different set of templates. We illustrate this common discrepancy in Appendix A. This inconsistency can lead to a misinterpretation of the reported results: the difference between the performance of two methods may be explained by the variation across prompt formats rather than the methods themselves.
In this work, we evaluate the template sensitivity of 19 models from 7 families, including state-of-the-art open-source models Touvron et al. (2023b); Almazrouei et al. (2023), and show that this issue persists irrespective of the model size and the number of demonstrations. Moreover, comparing various ICL enhancements while taking the template influence into account renders the superiority of one method over others less apparent. Therefore, it is likely that the gains reported for advanced prompting methods can often be attributed to a luckily chosen template.
Crucially, there are no universally best templates for a given task. The best performing demonstration format for a fixed evaluation setting (i.e., the dataset, the model, the demonstration set, and the prediction method) does not transfer consistently across models (even within the same family), demonstration sets, or different prediction methods. We find this concerning, as even the best template for a given setting can become produce poor results after slight changes, which makes “tuning” the template a very difficult task.
As a first step towards addressing template sensitivity in a practical way, we propose Template Ensembles — a test-time augmentation approach that averages model predictions over several task templates. This method is easy to implement and increases the average performance across templates for multiple prompting methods while reducing the sensitivity of these methods to the template choice.
@@
    We conduct a broad evaluation [^1] Our code and results of all evaluations can be found at github.com/yandex-research/mind-your-format of prompt template sensitivity across 19 models and 4 datasets, showing that the performance gains similar to using in-context learning improvements can be achieved solely by selecting a proper template.
@@ 2.1 In-Context Learning
An important property of LLMs is their ability to learn new tasks from only a few demonstrations Radford et al. (2019); Brown et al. (2020). This capability, known as in-context learning, forms the focus of our work. We focus on sequence classification, as it is the most widely studied task for understanding and improving ICL performance. Similarly, although instruction-finetuned models Sanh et al. (2022); Wei et al. (2021); Chung et al. (2022) are a popular choice for downstream adaptation in practice, we omit them from our analysis for consistency with prior work.
@@ 2.2 In-Context Learning Analysis
The choice of examples for ICL is highly important, as they enable the model to condition on correct input and label distributions for the task Wu et al. (2023); Nguyen and Wong (2023); Min et al. (2022d). Furthermore, the order of examples also significantly affects the results and does not transfer between models even within the same family Lu et al. (2022b); Zhao et al. (2021).
@@
Alternatively, one can use more advanced prediction methods that aim to reduce the variance across prompt formats. The Channel prompting technique, proposed in Min et al. (2022b), maximizes $P(x|y)$ instead of $P(y|x)$. The Calibration method Zhao et al. (2021) computes a correction factor based on the deviation of the model’s predictions for a placeholder input from a uniform distribution over labels and applies this factor to test set predictions.
However, as we show in Appendix A, these methods are evaluated on their own sets of templates. In this paper, we strive for a more unified view on the robustness of advanced prompting methods and compare their performance across a broader range of templates and models.
@@
Although the problem of prompt robustness is relatively well-known, until recently, the discussion of template robustness has been limited. Notably, Sclar et al. (2023) present a highly relevant study of prompt format sensitivity, reporting a significant performance variation across formats even for large models or minor template changes. The primary differences of our work are a simpler structure of the template space and a focus on the impact of template sensitivity on the evaluation of ICL modifications.
Moreover, several works study prompt robustness in a broader sense by considering models that use natural language instructions Webson and Pavlick (2022); Leidinger et al. (2023); Weber et al. (2023) instead of labeled demonstrations. Recently, Mizrahi et al. (2023) have shown that very similar instructions can lead to drastic differences in task performance for a variety of instruction-tuned models. Although we study a similar issue, we focus on in-context learning and the transfer of best templates between evaluation setups. Still, our findings agree with the results for instruction tuning, which confirms the necessity of language model evaluation that takes prompt design into account.
@@ 3.1 Models and Data
In preliminary experiments, we observed that the performance of some models in the few-shot regime lags behind their zero-shot results, indicating that the addition of context might, in fact, be more confusing for these models. Consequently, we excluded certain models from further investigation due to the high level of randomness in their behavior and for the sake of computational efficiency. Further details regarding the selection procedure can be found in Appendix B.
We experiment with 4 sequence classification datasets: SST-2 Socher et al. (2013), DBPedia ontology classification task Zhang et al. (2015a), AGNews Zhang et al. (2015b), and TREC Question Classification Li and Roth (2002). Although these datasets are frequently used in ICL studies, there is no consensus regarding the templates that should be used for each task.
@@ 3.2 Methods
Along with studying the robustness of standard in-context learning to template selection, we consider its improvements proposed in prior work. We consider two main directions of ICL enhancements mentioned in Section 2: example selection and prediction methods. For each setting, we aggregate the results over 3 random seeds for example selection, with 10 random templates used for each seed and report the mean and standard deviation of classification accuracy, unless specified otherwise.
As a baseline for demonstration selection, we choose the most straightforward approach of selecting $N$ random examples from the training dataset. Intuitively, selecting more relevant examples for ICL should yield better performance. In addition, we investigate the template sensitivity of two demonstration selection methods described in Section 2.1ITM and z-ICL. Specifically, we select examples for $N\in\{2,4\}$ using official implementations of each method .
Importantly, ITM requires training a concept model before choosing the examples. For GPT-2 Large, this procedure takes approximately 30 GPU hours on a single NVIDIA A100-80GB. Repeating it for each of the evaluated models would be infeasible, especially given that the largest model has 86 times more parameters than GPT-2 Large. Therefore, we use the checkpoints of the GPT2-Large concept model provided by the authors to select demonstrations for all datasets. Also, we reuse the same set of examples for all models, leveraging authors’ observations that demonstrations chosen with ITM can be transferred between models.
As discussed in Section 2, more advanced prediction techniques can improve in-context learning accuracy. Therefore, we compare Direct with Channel Min et al. (2022b) and Calibration Zhao et al. (2021) prompting methods.
@@ 4.1 Baseline Results
We begin with analyzing the robustness of language models to the template choice in the baseline setup. Specifically, we evaluate models in zero-shot and few-shot settings, selecting 2/4 random demonstrations and using the Direct prediction method. Our results in Table 3 show that even the most advanced models such as Falcon and LLaMA 2 are highly sensitive to the prompt formatting; Appendix C contains the results for the full set of 19 models. This sensitivity makes it harder to observe the increase in ICL performance with the model size or the number of demonstrations, but this trend still persists. However, even the largest models exhibit the standard deviation reaching 35% of their respective scores.
To mitigate this lack of robustness, we could remove consistently underperforming prompt formats from our template pool. We analyze the impact of separate template components in Appendix D and find that there are no particular parts (e.g, verbalizers or separators) which could be excluded from evaluation. Furthermore, we observe that a combination of “suboptimal” parts may result in an optimal template.
@@
Table 3: Classification accuracy in the baseline setting for 2 LLMs families. Standard deviation across 30 runs (10 templates for 3 sets of demonstrations) is in underscript. The report for all models is presented in Appendix C.
@@ 4.2 Prediction Methods
We evaluate Channel and Calibration methods in a 2-shot setting along with the Direct baseline.
Figure 2: Comparison of the in-context learning prediction methods in the 2-shot setting.
As depicted in Figure 2, both Channel and Calibration exhibit improved performance as compared to Direct. Still, for a range of models and datasets, the range of scores for Direct substantially overlaps with those of both advanced methods. This indicates the existence of templates that reach the best performance with the Direct prediction method.
Additionally, Table 10 of  Appendix E reveals that despite Calibration yields the highest mean accuracy most often, it is much more susceptible to template choice than Channel. Therefore, there is no discernible pattern to select between these two methods for a given dataset and model pair.
@@ 4.3 Example Selection Methods
Another area of ICL improvements that we evaluate on the matter of template sensitivity is the example selection strategy. We compare ITM and z-ICL methods to Random baseline in 4-shot setting since it was the main evaluation setting in the works proposing these methods. We use Direct prediction method to evaluate the gains of advanced example selection strategies independently from other concurrent modifications.
Results in Figure 3 and Table 12 illustrate that when considering template sensitivity issue, advanced example selection methods often perform worse than random choice baseline. ITM method increases the average performance in most cases but still has a remarkably high standard deviation across templates. Examples selected using z-ICL method lead to consistent but poor performance.
Figure 3: Comparison of the selection methods in the Direct 4-shot setting. For the evaluation results of other models and datasets please refer to Appendix F.
Note that our evaluation setup differs from those in the works proposing these methods. Namely, we use Direct prediction method and sample 10 random templates that might have not included the templates used by authors of ITM and z-ICL.
We conclude from this evaluation that a prompt format selection is as important axis of ICL enhancements as the choice of a proper selection or prediction method. However, unlike the other directions for improving ICL, the search space of possible templates is infinite, making exhaustive exploration for each combination of dataset, model and number of examples impractical. Ideally, the best template for one setting would be optimal for others, or at least for similar settings. However, this is not the case.
@@ 5 Template Transfer Evaluation
5.1 Transfer Between Setups
We begin with defining what is considered a successful transfer between setups. In order to do so, we evaluate how the quality of models varies across 30 random templates from Section 3.1. The results detailed in Appendix G demonstrate the top-10 template on average yields 0.90 ± 0.04 of the best template score. Therefore, if a prompt format is present in top-10 for both compared setups, we consider this template successfully transferred.
To compare sets of the best templates for a pair of settings $A,B$ we employ an Intersection over Union (IoU), also known as a Jaccard similarity coefficient  Jaccard (1912), for the top-10 templates: IoU(A,B) = |A ∩B||A ∪B| = |A ∩B||A| + |B| - |A ∩B|
Another measure of transferability could be a Spearman’s rank correlation coefficient Zar (2005). However, it is vulnerable to the case when the worst templates share similar ranking; it boosts the score, while transfer of efficient templates remains low. We illustrate this issue in Appendix H.
@@ 5.2 Transfer Between Models
Next, we analyze transfer of the best performing templates between models in the baseline setup. Specifically, we collect the results of each model in 2-shot with Direct prediction and Random selection methods for 30 templates. A heatmap of the IoU for the top-10 best templates between 19 models on the DBPedia dataset is presented in Figure 4. For other datasets, please see Appendix I.
We observe that the IoU exceeds 0.5 only for few model pairs on all datasets, signifying that generally template transferability between models in the same setup is rather low. This is especially surprising for models within the same family.
These observations demonstrate that, given the sensitivity to prompt format, it is inappropriate to compare methods’ performance across models using a single template, since a template that is effective for one model can easily turn out to be one of the worst choices for another model.
Figure 4: IoU of top-10 templates for all models in Random 2-shot with Direct prediction method on the DBPedia dataset.

### v3

@@ Abstract
Large language models demonstrate a remarkable capability for learning to solve new tasks from a few examples. The prompt template, or the way the input examples are formatted to obtain the prompt, is an important yet often overlooked aspect of in-context learning. In this work, we conduct a comprehensive study of the template format’s influence on the in-context learning performance. We evaluate the impact of the prompt template across 21 models (from 770M to 70B parameters) and 4 standard classification datasets. We show that a poor choice of the template can reduce the performance of the strongest models and inference methods to a random guess level. More importantly, the best templates do not transfer between different setups and even between models of the same family. Our findings show that the currently prevalent approach to evaluation, which ignores template selection, may give misleading results due to different templates in different works. As a first step towards mitigating this issue, we propose Template Ensembles that aggregate model predictions across several templates. This simple test-time augmentation boosts average performance while being robust to the choice of random set of templates.
Mind Your Format: Towards Consistent Evaluation
of In-Context Learning Improvements
Anton Voronov Yandex, HSE University, MIPT voronov.ad@phystech.edu                        Lena Wolf Yandex, HSE University e.a.volf@yandex.ru                        Max Ryabinin Together AI mryabinin0@gmail.com
@@ 1 Introduction
These few-shot or in-context learning (ICL) abilities Brown et al. (2020) of large models are a subject of frequent study, as the primary factors behind them are not yet fully understood. For example, one line of work investigates in-context learning within different theoretical frameworks Xie et al. (2022); Garg et al. (2022); Akyürek et al. (2023). In addition, multiple publications study the importance of different prompt attributes, such as the order of input demonstrations (Lu et al., 2022a) and their labels (Min et al., 2022d).
@@
As shown in Zhao et al. (2021); Min et al. (2022a), the prompt format (i.e., a transformation from a set of examples to a natural language input) is also highly important. However, this aspect is often overlooked in most existing studies. Namely, works proposing modifications of ICL frequently present their results for a specific template without specifying the criteria guiding its selection. Furthermore, even when the results are averaged over a set of templates, they are compared to methods that were evaluated on a different set of templates. We illustrate this common discrepancy in Appendix A. Such inconsistency can lead to a misinterpretation of the reported results: the difference between the performance of two methods may be explained by the variation across prompt formats rather than the methods themselves.
In this work, we evaluate the template sensitivity of 21 models from 8 families, including state-of-the-art open-access models, such as Llama 2 Touvron et al. (2023b) and Falcon Almazrouei et al. (2023), as well as latest instruction-tuned models, such as Mistral Jiang et al. (2023) and Llama 3 Instruct AI@Meta (2024). We show that this issue persists regardless of the model size and the number of demonstrations. Moreover, comparing various in-context learning enhancements while taking the template influence into account renders the superiority of one method over others less apparent. Therefore, it is likely that the gains reported for advanced prompting methods can often be attributed to a luckily chosen template.
Crucially, there are no universally best templates for a given task. The best performing demonstration format for a fixed evaluation setting (i.e., the dataset, the model, the demonstration set, and the prediction method) does not transfer consistently across models (even within the same family), demonstration sets, or different prediction methods. We find this concerning, as even the best template for a given setting can produce poor results after slight changes, which makes “tuning” the template a very difficult task.
As a first step towards addressing template sensitivity in a practical way, we propose Template Ensembles — a test-time augmentation approach that averages model predictions over several prompt formats. This method is easy to implement and increases the average performance across templates for multiple prompting methods while reducing the sensitivity of these methods to the template choice.
@@
    We conduct a broad evaluation [^1] Our code and results of all evaluations can be found at github.com/yandex-research/mind-your-format of prompt template sensitivity across 21 models and 4 datasets, showing that the performance gains similar to using in-context learning improvements can be achieved solely by selecting a proper template.
@@ 2.1 In-Context Learning
An important property of LLMs is their ability to learn new tasks from only a few demonstrations Radford et al. (2019); Brown et al. (2020). This capability, known as in-context learning, forms the focus of our work. We focus on sequence classification, as it is the most widely studied task for understanding and improving ICL performance.
@@ 2.2 In-Context Learning Analysis
The choice of examples for ICL is highly important, as they enable the model to condition on correct input and label distributions for the task Wu et al. (2023); Nguyen and Wong (2023); Min et al. (2022d); Chang and Jia (2023). Furthermore, the order of examples also significantly affects the results and does not transfer between models even within the same family Lu et al. (2022b); Zhao et al. (2021).
@@
Alternatively, one can use more advanced prediction methods that aim to reduce the variance across prompt formats. The Calibration method Zhao et al. (2021) computes a correction factor based on the deviation of the model’s predictions for a placeholder input from a uniform distribution over labels and applies this factor to test set predictions. Recent work has proposed multiple improvements of this method Fei et al. (2023); Han et al. (2023); Zhou et al. (2024); to limit the scope of our study, we focus only on the base Calibration approach in this work. Lastly, the Channel prompting technique, proposed in Min et al. (2022b), maximizes $P(x|y)$ instead of $P(y|x)$.
Both of these methods aim to mitigate the issue of ICL sensitivity to the prompt template choice. However, as we show in Appendix A, these methods are evaluated on their own sets of templates. In this paper, we strive for a more unified view on the robustness of advanced prompting methods and compare their performance across a broader range of templates and models.
@@
Although the problem of prompt robustness is relatively well-known, until recently, the discussion of template robustness has been limited. Notably, Sclar et al. (2023) present a highly relevant study of prompt format sensitivity, reporting a significant performance variation across formats even for large models or minor template changes. While their experiments are conducted in the standard setup (randomly selected examples and default prompting), our work instead focuses on alternative prompting and example selection methods, several of which Zhao et al. (2021); Min et al. (2022b) were proposed to improve the prompt robustness of ICL. Similarly to papers in other subfields of machine learning arguing for a more consistent methodology Dacrema et al. (2019); Musgrave et al. (2020); Platonov et al. (2023), the goal of our work is to demonstrate that disparate experiment setups lead to an invalid comparison of competing methods.
Moreover, several works study prompt robustness in a broader sense by considering models that use natural language instructions instead of labeled demonstrations Webson and Pavlick (2022); Leidinger et al. (2023); Weber et al. (2023). Recently, Mizrahi et al. (2023) have shown that very similar instructions can lead to drastic differences in task performance for a variety of instruction-tuned models. Although we study a similar issue, the focus of our work is on in-context learning and the transfer of best prompts between evaluation setups. Still, we find that instruction-tuned models lack in-context robustness as well, which confirms previous observations and emphasizes the need for language model evaluation that takes prompt design into account.
@@ 3 Setup & Methodology
Llama 3 Instruct AI@Meta (2024) 8
Mistral v0.3 Instruct Jiang et al. (2023) 7
@@ 3.1 Models and Data
In preliminary experiments, we observed that the performance of some models in the few-shot regime lags behind their zero-shot results. Hence, we excluded these models from further investigation. Further details regarding this selection procedure can be found in Appendix B.
We experiment with 4 sequence classification datasets: SST-2 Socher et al. (2013), DBPedia ontology classification task Lehmann et al. (2015), AGNews Zhang et al. (2015), and TREC Question Classification Li and Roth (2002). Although these datasets are frequently used in ICL studies, there is no consensus regarding the templates that should be used for each task.
Dataset Input verbalizer Output verbalizer Intra-separator Inter-separator
SST-2 “output: {}”, “target: {}”, “label: {}”,
[…]
“It was {}.”, “All in all {}.”, “A {} piece.”
DBPedia “output: {}”, “target: {}”, “label: {}”,
AGNews “Topic: {}.”, “Subject: {}.”,
TREC ‘This is about {}.”, “It is about {}.”
@@ 3.2 Methods
Along with studying the robustness of standard in-context learning, we consider its improvements proposed in prior work. We focus on two main directions of ICL enhancements mentioned in Section 2: example selection and prediction methods. For each setting, we aggregate the results over 3 random seeds for example selection, with 10 random templates used for each seed and report the mean and standard deviation of classification accuracy, unless specified otherwise.
As a baseline for demonstration selection, we choose the most straightforward approach of selecting $N$ random examples from the training dataset. Intuitively, selecting more relevant examples for ICL should yield better performance. Therefore, we investigate the template sensitivity of two demonstration selection methods described in Section 2.1ITM and z-ICL. Specifically, we select $N=4$ examples using official implementations of each method.
Importantly, ITM requires training a concept model before choosing the examples. For GPT-2 Large, this procedure takes approximately 30 hours on a single NVIDIA A100-80GB. Repeating it for each model would be infeasible, especially given that the largest model has 86 times more parameters. Therefore, we use the checkpoints of the GPT2 Large concept model provided by the authors to select demonstrations. Also, we reuse the same examples for all models, leveraging authors’ observations that demonstrations chosen with ITM can be transferred between models.
As discussed in Section 2, more advanced prediction techniques can improve in-context learning accuracy. Therefore, we compare Direct prompting with Channel Min et al. (2022b) and Calibration Zhao et al. (2021) prediction methods.
@@ 4.1 Baseline Results
We begin with analyzing the robustness of language models to the template choice in the baseline setup. Specifically, we evaluate models in zero-shot and few-shot settings, selecting 2/4 random demonstrations and using the Direct prediction method.
Our results in Table 3 show that even the most capable models such as Falcon and Llama 2 are highly sensitive to the prompt format; Appendix C contains the results for the full set of 19 base models, and Appendix L reports the results for instruction-tuned models. Although the variance caused by this sensitivity makes it harder to observe the increase in ICL performance with the model size or the number of demonstrations, both trends still persist. However, even the largest models have standard deviations of scores up to 35% of their mean values.
To mitigate this lack of robustness, we could remove consistently underperforming prompt formats from the template pool. We analyze the impact of separate template components in Appendix D and find that there are no specific parts (for example, verbalizers or separators) which could be excluded from evaluation. Furthermore, we observe that a combination of “suboptimal” parts may result in an optimal template.
@@
Table 3: Classification accuracy in the baseline setting for 2 LLM families. Standard deviation across 30 runs (10 templates for 3 sets of demonstrations) is in underscript. The results for all base models are presented in Appendix C.
@@ 4.2 Prediction Methods
We evaluate Channel and Calibration methods in the 2-shot setting along with the Direct baseline.
Figure 2: Comparison of in-context learning prediction methods in the 2-shot setting.
As depicted in Figure 2, both Channel and Calibration generally exhibit improved performance in comparison with Direct. Still, for a number of models and datasets, the range of scores for Direct substantially overlaps with those of advanced methods. This suggests that there are templates reaching the best performance with the Direct prediction method.
Additionally, Table 10 of Appendix E reveals that despite Calibration yielding the highest mean accuracy more often than other methods, it is more sensitive to the template choice than Channel. Similar findings for instruction-tuned models are contained in Appendix L. Therefore, the choice of the prediction method should likely rely on the downstream usage scenario and the target evaluation setting.
@@ 4.3 Example Selection Methods
Another area of ICL improvements that we evaluate on the matter of template sensitivity is the example selection strategy. We compare ITM and z-ICL methods to the Random baseline in 4-shot setting, since using 4 demonstration was the main evaluation setting in the works proposing these methods. We use Direct prediction method to evaluate the gains of advanced example selection strategies independently from other ICL modifications.
Results in Figure 3 and Table 14 illustrate that when taking template sensitivity into account, advanced example selection methods often perform worse than random choice baseline. ITM increases the average performance in most cases but still has a remarkably high standard deviation across templates. Examples selected using the z-ICL method lead to more consistent but worse performance.
Figure 3: Comparison of the selection methods in the Direct 4-shot setting. For the evaluation results of other models and datasets, please refer to Appendix G.
Note that our evaluation setup differs from those described in the original works, which might explain the discrepancy between our findings and the results reported by authors. Namely, we use the Direct prompting method and sample 10 random templates that may not include the templates used by authors of ITM and z-ICL. To confirm template instability for prediction methods in their original implementations, we reproduce both methods and report our findings in Appendix F. We observe high sensitivity to the prompt format, which raises a question of how much the reported gains of these methods can be attributed to the methods themselves and not to the template choice.
We conclude that the prompt format should be viewed as important as the example selection or the prediction method for ICL evaluation. However, the search space of possible templates is infinite, which makes exhaustive search for each combination of the dataset, the model and the number of examples impractical. Ideally, the best template for one setting would be optimal for all others or at least for similar settings. However, as we demonstrate in the following section, this is not the case.
@@ 5 Template Transfer Evaluation
5.1 Setup
We begin by defining a successful transfer between ICL settings. In order to do so, we evaluate how the quality of model predictions varies across 30 random templates from Section 3.1. The results described in Appendix H demonstrate that the top-10 template on average yields 90% of the best template score. Therefore, if a prompt format is present in top-10 for both of the two compared setups, we can consider this an instance of successful transfer.
To compare sets of the best templates for a pair of settings, we compute Intersection-over-Union (IoU), also known as the Jaccard similarity coefficient  Jaccard (1912), for top-10 best templates in each setting. We also consider using the $\rho$ rank correlation coefficient Spearman (1904) as another measure of template transfer. However, its value can increase when low-performing templates have similar rankings in different ICL setups, while the transfer of efficient templates remains low. Still, we provide the results for this metric in Appendix I.
@@ 5.2 Transfer Between Models
Next, we analyze the transfer of the best-performing templates between models in the baseline setup. Specifically, we collect the results of each model in the 2-shot learning setting with Direct prediction method and Random demonstrations (fixed throughout the experiments) for 30 templates. A heatmap of IoU for the transfer of top-10 best templates between 19 base models on the DBPedia dataset is presented in Figure 4; for other datasets, please see Appendix J.
We observe that the IoU values exceed 0.5 only for a few model pairs on all datasets, meaning that the capacity for template transfer between models in the same setup is generally low. This is especially concerning for models within a single family: as these models are trained on the same data and have the same architecture, one would expect them to perform similarly on the same prompt formats.
These observations signify that comparing ICL methods across models with a single template can lead to incorrect conclusions: a template that is effective for one model can easily be one of the worst choices for another model.
Figure 4: IoU of top-10 templates for all base models with 2 random demonstrations and the Direct prediction method on the DBPedia dataset.